OpenStudy (anonymous):

what is 102 divided by 2 in base 3

OpenStudy (anonymous):

base 3 or mod 3?

OpenStudy (anonymous):

base3

OpenStudy (anonymous):

sorry for the delay, my computer's lagging. So in base 3, 102 means our usual 11, is that correct?

OpenStudy (anonymous):

no in base three you have a one and a two three would be represented by 10 or one zero 4 is 11 or one one and so on

OpenStudy (anonymous):

So 102 = (3^2) + 0* (3^1) + 2*(3^0), wouldnt it?

OpenStudy (anonymous):

base 3 means 2 + 0*3 + 2*3*3 = 2+2*9 = 2+18 = 20 in base 10, right?

OpenStudy (anonymous):

Then you just divide by 2, and convert back to base 3.

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

Well, base 10 (which is what we use) has one digit in each place. A digit in the 10^0 (1's) place. A digit in the 10^1 (10's) place. A digit in the 10^2 (100's) place, and so on. base 3 is the same way.

OpenStudy (anonymous):

A digit in the 3^0 (1's) place. A digit in the 3^1 (3's) place. A digit in the 3^2 (9's) place.

OpenStudy (anonymous):

So reading 102 in base 10 is what you're familiar with. That would be 1 100 + 2 1's.

OpenStudy (anonymous):

Just take that and apply it to three. Looking at it in that way, can you say what 102 in base 3 would be?

OpenStudy (anonymous):

Isn't it 11?

OpenStudy (anonymous):

2 + 0*3 + 2*3*3 would be 202

OpenStudy (anonymous):

Yea, misread the question earlier. Changed it to what it was supposed to be without realizing.

OpenStudy (anonymous):

Just noticed I did that and I apologize.

OpenStudy (anonymous):

thank you for all your help

OpenStudy (anonymous):

Yeah, either there's a typo in the book, or they want you to do decimals in base 3. 11/2 = 5.5 = 5 and 1/2. 1/2 = 1/3 + 1/9 + 1/27 + ... = 0.111111... and 5 is written 12, so the answer would be 12.1111...